OpenStudy (anonymous):

What is the exact value of 7q^4/√q^15

OpenStudy (anonymous):

Here, you'd like to get an even degree of q under the radical so that you can pull it out. What is the next even number smaller than 15?

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

Yes, so now you have \[7q^4 \sqrt{q^{14} q}\] Right? So you can take the q^14 out of the radical. Remember that sqrt(q^2) = q, when you pull it out.

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

When you pull q^14 out of the radical, you will have q^7, not q^14.

OpenStudy (anonymous):

So now what do you have?

OpenStudy (anonymous):

I did not see the divide, but the same still applies here, but you have: \[7q^4 / \sqrt{q^{14}q}\]

OpenStudy (anonymous):

\[\frac{7q^4}{\sqrt{q^{15}}}\]